jadadunn:

Solve this system of equations algebraically: y – 10 = 11x + x2 y – 12x = 30 The first solution is (–4, –18). The second solution is

jadadunn:

@JustSaiyan

JustSaiyan:

@Vocaloid

vaporeon246:

This might help you https://www.mathway.com/Algebra

Vocaloid:

y – 10 = 11x + x2 y – 12x = 30 you can re-write these both in terms of y to get y = 10 + 11x + x^2 y = 12x + 30 then set these equations equal to each other to get 10 + 11x + x^2 = 12x + 30 move everything to one side and then apply the quadratic formula